please help me solve this: the area and the perimeter have to be the same on a rectangle what could they be?
Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

Area = xy
Perimeter = 2(x+y)
------------
2x + 2y = xy
The solution can be any point on the hyperbola where both x and y are positive.
Also (0,0) is a solution.

Answer by checkley77(12844)

Area=perimeter.
xy=2x+2y
xy=2(x+y)
3*6=2(3+6)
18=2*9
18=18